if the period of a simple pendulum is 2 s, what should its length be ? if the length of the pendulum is doubled what will its period be.


1
Expert's answer
2022-02-28T10:41:17-0500
"T_0=2\pi\sqrt{l_0/g}"

"l_0=gT_0^2/(4\pi^2)=9.8*2^2/(4*3.14^2)=1.0\:\rm m"

"T=2\pi\sqrt{2l_0/g}=\sqrt{2}T_0=2.8\:\rm s"